# Maximal discs of Weil-Petersson class in $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$

### v1

@@ Appendix A Anti de Sitter space of dimension 3
Let $\mathbb{R}^{2,2}$ denote the pseudo-Euclidean 4-space endowed with linear coordinates $\mathbf{x}=(x_{1},x_{2},x_{3},x_{4})$. Let $q$ denote the quadratic form defined by
@@
One can check that $\mathbb{H}^{2,1}$ is a smooth connected submanifold of $\mathbb{R}^{2,2}$ of dimension $3$. The tangent space $T_{\mathbf{x}}\mathbb{H}^{2,1}$ can be identified with a subspace of $\mathbb{R}^{2,2}$, which is given by
@@
Let us define the Anti-de Sitter space of dimension 3 by
(A.3) $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}:=\mathbb{H}^{2,1}/\{\pm\mathrm{Id}\}$
where $\mathrm{Id}$ denotes the identity element in $\mathrm{O}(2,2)$. Then $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ is a Lorentzian manifold with the induced metric from $\mathbb{H}^{2,1}$ and has the constant curvature $-1$. By the definition, $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ is identified with a subset of real projective space $\mathbb{RP}^{3}$, that is
@@
Let $\mathrm{M}(2,\mathbb{R})$ denote the vector space of $2\times 2$ matrices with real entries. There is an isometric identification between $(\mathrm{M}(2,\mathbb{R}),-\mathrm{det})$ and $(\mathbb{R}^{2,2},q)$. Under this isomorphism, $\mathbb{H}^{2,1}$ is identified with the Lie group $\mathrm{SL}(2,\mathbb{R})$. The Lie group $\mathrm{SL}(2,\mathbb{R})$ has the Killing form $\kappa$, that is a bi-invariant bilinear form on the Lie algebra $\mathfrak{sl}(2,\mathbb{R})$ of the signature $(2,1)$. The bilinear form $\kappa$ induces a Lorentzian metric on $\mathrm{SL}(2,\mathbb{R})$, which we denote by $g_{\kappa}$. Recalling (A.3), one can check the Anti-de Sitter space $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ is the same as $\mathrm{PSL}(2,\mathbb{R})$ equipped with $\frac{1}{8}g_{\kappa}$ as a Lorentzian manifold.
Let us notice that $\mathrm{SL}(2,\mathbb{R})\times\mathrm{SL}(2,\mathbb{R})$ acts on $\mathrm{M}(2,\mathbb{R})$ by
$(\alpha,\beta)\cdot\gamma=\alpha\circ\gamma\circ\beta^{-1}\qquad\text{for}\quad\gamma\in\mathrm{M}(2,\mathbb{R}),$
and this action preserves the quadratic form $-\mathrm{det}\cong q$. Hence, we have
@@
where $K:=\{(\mathrm{Id},\mathrm{Id}),(-\mathrm{Id},-\mathrm{Id})\}$, so that
@@
By (A.4), for the projectivization of $\mathrm{M}(2,\mathbb{R})$ denoted by $\mathrm{P}(\mathrm{M}(2,\mathbb{R}))$,
@@
We have the following homeomorphism
(A.6) $\partial\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}\,\to\,\mathbb{R}\mathbb{P}^{1}\times\mathbb{R}\mathbb{P}^{1}$
defined by
@@
A timelike geodesic in $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ has the following form
$L_{p,q}=\{\gamma\in\mathrm{PSL}(2,\mathbb{R})\,|\,\gamma(q)=p\,\}$
@@
is defined by $G(x)=(p,q)$ where $L_{p,q}$ is the timelike geodesic orthogonal to the tangent space of $\Sigma$ at $x$.

### v2

@@ Appendix A Anti de Sitter space of dimension 3
Let $\mathbb{R}^{2,2}$ denote the pseudo-Euclidean 4-space with linear coordinates $\mathbf{x}=(x_{1},x_{2},x_{3},x_{4})$. Consider the quadratic form
@@
One can verify that $\mathbb{H}^{2,1}$ is a smooth connected $3$-dimensional submanifold of $\mathbb{R}^{2,2}$. The tangent space $T_{\mathbf{x}}\mathbb{H}^{2,1}$ at $x\in\mathbb{H}^{2,1}$ is identified with the subspace
@@
The $3$-dimensional Anti-de Sitter space is then defined as
(A.3) $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}:=\mathbb{H}^{2,1}/\{\pm\mathrm{Id}\},$
where $\mathrm{Id}$ is the identity element in $\mathrm{O}(2,2)$. The space $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ inherits the Lorentzian metric from $\mathbb{H}^{2,1}$ and has the constant curvature $-1$. By the definition, $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ can be identified with a subset of the real projective space $\mathbb{RP}^{3}$:
@@
Let $\mathrm{M}(2,\mathbb{R})$ denote the vector space of $2\times 2$ real matrices. There is an isometric identification between $(\mathrm{M}(2,\mathbb{R}),-\mathrm{det})$ and $(\mathbb{R}^{2,2},q)$, under which the hyperboloid $\mathbb{H}^{2,1}$ corresponds to the special linear group $\mathrm{SL}(2,\mathbb{R})$. The Lie group $\mathrm{SL}(2,\mathbb{R})$ has a bi-invariant bilinear form, known as the Killing form $\kappa$, on its Lie algebra $\mathfrak{sl}(2,\mathbb{R})$. The Killing form $\kappa$ has the Lorentzian signature $(2,1)$, inducing a Lorentzian metric on $\mathrm{SL}(2,\mathbb{R})$, which we denote by $g_{\kappa}$. From (A.3), one can verify that the Anti-de Sitter space $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ is naturally identified with $\mathrm{PSL}(2,\mathbb{R})$ equipped with the Lorentzian metric $\frac{1}{8}g_{\kappa}$.
The group $\mathrm{SL}(2,\mathbb{R})\times\mathrm{SL}(2,\mathbb{R})$ acts on $\mathrm{M}(2,\mathbb{R})$ via
$(\alpha,\beta)\cdot\gamma=\alpha\circ\gamma\circ\beta^{-1}\qquad\text{for}\quad\gamma\in\mathrm{M}(2,\mathbb{R}).$
This action preserves the quadratic form $-\mathrm{det}\cong q$, leading to the identification
@@
where $K:=\{(\mathrm{Id},\mathrm{Id}),(-\mathrm{Id},-\mathrm{Id})\}$. Consequently, the connected component of the isometry group of $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ is given by
@@
By (A.4), the boundary of $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ in the projectivized space $\mathrm{P}(\mathrm{M}(2,\mathbb{R}))$ is given by
@@
This boundary admits the following homeomorphism:
(A.6) $\partial\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}\,\to\,\mathbb{R}\mathrm{P}^{1}\times\mathbb{R}\mathrm{P}^{1},$
which is explicitly defined by
@@
Timelike geodesics in $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ are given by
$L_{p,q}=\{\gamma\in\mathrm{PSL}(2,\mathrm{R})\,|\,\gamma(q)=p\,\}$
@@
is defined by
$G(x)=(p,q),$
where $L_{p,q}$ is the timelike geodesic orthogonal to the tangent space of $\Sigma$ at $x$.